OpenStudy (anonymous):

If the pool is 25 feet long and 20 feet wide, find the area of the pool and deck.

OpenStudy (paki):

area of pool= 25*20= 500

OpenStudy (anonymous):

What about the deck then?

OpenStudy (anonymous):

Do you know how to do it in FOIL? For example: A contractor is building a deck around a rectangular swimming pool. The deck is x feet from every side of the pool. Write an expression for the total area of the pool and deck. Area = length · width = (2x + 20)(2x + 15) = (2x)(2x) + (2x)(15) + (20)(2x) + (20)(15) =4x2 +30x+40x+300 =4x2 +70x+300 So, the total area of the deck and pool is 4x2 + 70x + 300.